# GenView: Enhancing View Quality with Pretrained Generative Model for Self-Supervised Learning

### v1

@@ 1 Introduction
Self-supervised learning (SSL) has demonstrated remarkable capability in acquiring robust and generalized visual representations from abundant unlabeled data sources [62, 26, 92, 63, 33, 15, 29, 102, 38, 34, 50, 2, 25], which can be transferred or leveraged in downstream tasks. Among the various approaches within SSL, Contrastive Learning (CL) [15, 17, 12, 13, 18, 102, 38, 51] has emerged as a prominent method, showcasing its effectiveness in numerous downstream tasks (e.g., classification [31, 48, 96], detection [71, 47, 32, 9, 86], and segmentation [27, 58, 90, 49, 52]). CL aims to learn invariant representations that remain consistent across various conditions or environments by maximizing the similarity of representations obtained from different distorted versions of a sample, referred to as positive views. Consequently, the construction of high-quality positive views is crucial for CL. A high-quality positive view should retain the semantics of the original images while introducing as much semantic-irrelevant attribute diversity and environmental variations as possible, such that the learned representations can be more generalizable for downstream tasks.
@@ 2 Related Work
2.1 Self-Supervised Learning
@@ 2.2 Generative Models
Instead of training generative models from scratch, some methods use pretrained generative models to augment representation learning, leveraging the prior knowledge learned from large-scale datasets [77, 14] to enhance the high-level diversity of the generated views [40, 35, 83, 21, 79, 84, 100]. However, these models rely on constant [35, 21, 83, 100] or random [79, 84, 103] hyperparameters to determine the extent of deviation in the generated images. This can lead to uncontrolled data generation characterized by inconsistent semantics with the conditional image, reducing the quality of positive pairs. In contrast, our approach employs an adaptive view generation that controls the noise level when sampling images to keep a balance between semantic fidelity and diversity based on individual image characteristics. We also propose a quality-driven contrastive loss to enhance the contributions of high-quality positive pairs while diminishing the impact of low-quality and false pairs.
@@ 3.4 Quality-driven Contrastive Loss
$\mathbf{z}^{b}_{i}=\mathbf{M}_{i}^{b}\otimes\mathbf{F}_{i},\quad\mathbf{z}^{b+}_{i}=\mathbf{M}_{i}^{b+}\otimes\mathbf{F}_{i}^{+},$ (11)
@@ 4 Experiments
We compare GenView with state-of-the-art SSL methods, including MoCov2 [33], BYOL [29], SwAV [12], SimSiam [16], and MoCov3 [18]. We experiment with various network architectures, such as ResNet-18 [31], ResNet-50 [31], ViT-S [20], and ViT-B [20]. By default, ResNet-50 serves as the backbone. ViT-S and ViT-B are adopted for comparison with MoCov3. For details on adaptive view generation and quality-driven contrastive loss implementations for different pretraining datasets, please refer to the Appendix A.
@@ 4.1 Main Results
Linear classification. GenView is framework-agnostic, allowing flexibility with SSL frameworks and associated training components like backbone networks, loss functions, and optimizers. To ensure fair comparisons, we maintain consistent pretraining settings as baseline methods on ImageNet-1K [19] (IN-1K). To evaluate our method, we follow a standard linear classification protocol, as described in previous works [15, 17, 29]. The linear classifier is trained on top of the frozen representation for 90 epochs with a batch size of 1,024, an initial learning rate of 0.4, an SGD optimizer with 0.9 momentum and no weight decay, and the cosine-annealed learning rate schedule [59]. For ViT-based models, the initial learning rate is set to 12. Tab. 1 presents the results of top-1 accuracy on the validation set of IN-1K. GenView consistently improves SSL performance across various frameworks, including ResNet-50 and Transformer architectures like ViT-S and ViT-B. Its effectiveness is maintained across different pretraining epochs, outperforming the MoCov3 baselines pretrained for 100 or 300 epochs. GenView outperforms C-Crop [65] that also constructs better views, highlighting our advantage in utilizing pretrained generative models’ prior knowledge to create diverse views in a controlled manner. GenView can complement both contrastive (e.g. MoCov2 and MoCov3) and non-contrastive methods (e.g. BYOL, SimSiam, and SwAV), addressing their limitations of positive pair quality. These results demonstrate GenView’s consistent ability in enhancing the linear classification performance of various SSL models. It’s noted that when GenView is integrated with MoCov3 utilizing a ResNet-50 backbone and pretrained over 300 epochs, it achieves competitive performance (74.8% with 1.28 million images) compared to CLIP (74.3% on WebImageText with 400 million pairs), highlighting GenView’s efficiency.
Table 1: Comparison with existing SSL methods under linear evaluation on IN-1K. ***: our reproduction.
@@
Semi-supervised classification. We evaluate the fine-tuning performance of the pretraind models for semi-supervised classification with 1% and 10% of labeled IN-1K samples, selected by SimCLR [15]. We fine-tune the models for 20 epochs with the classifier learning rate 1.0 (0.2) and backbone learning rate 0.00001 (0.02) for 1% (10%) subset with a cosine-annealed scheduler. Tab. 2 presents the results of top-1 and top-5 accuracy on the validation set of IN-1K. Our method consistently outperforms the baseline approaches across different training durations. With 1% labels, GenView pretrained for 200 epochs with MoCov2 achieves an improvement of +8.5% in top-1 accuracy, and the one pretrained for 300 epochs with MoCov3 still improves top-1 accuracy by +1.9%.
Table 2: Comparison with existing SSL methods for semi-supervised learning on IN-1K. Models with ResNet-50 backbone are pretrained on IN-1K. ***: our reproduction.
@@
Transfer learning on object detection and instance segmentation.
Table 3: Transfer learning on MS-COCO object detection and instance segmentation. Models with ResNet-50 backbone are pretrained for 200 epochs on IN-1K. ***: our reproduction.
@@
Comparison with other view construction methods.
@@ 4.2 Ablations
Influence of each component. We evaluate the contributions of individual components as well as their combinations. ResNet-18 models are pretrained on IN-100 for 100 epochs using MoCov3 as the baseline, with a batch size of 512. IN-100 is a subset of IN-1K selected by [81]. For conditioning the generation of positive views with GenView, we employ 50,000 randomly selected class-balanced images from IN-100. We use a cosine decay learning rate schedule and employ the LARS optimizer with a learning rate of 1.2, weight decay of 1e-6, and momentum of 0.9. Linear evaluation settings are consistent with those detailed in Tab. 1, with a training duration of 50 epochs.
@@
$\checkmark$ $\checkmark$ $\checkmark$ 75.40 (normal-$\uparrow$ 9.88)
Tab. 6 offers valuable insights: (1) Utilizing our framework but without our adaptive view generation significantly enhances accuracy, achieving a top-1 accuracy improvement of 5.98% compared to the baseline. (2) The incorporation of adaptive view generation further elevates model performance, resulting in an improvement of 8.44% (from 65.52% to 73.96%). (3) The quality-driven contrastive loss also plays a pivotal role in our framework. It can further improve the performance of adaptive view generation. Applying the quality-driven contrastive loss to the baseline method leads to a modest gain of 1.45% (from 65.52% to 66.97%). However, when combined with our framework, a more substantial performance improvement of 3.38% (from 71.50% to 74.88%) is observed. This highlights the effectiveness of our framework and also the importance of the proposed modules in enhancing contrastive learning by improving the quality of positive pairs.
Influence of the noise level selection strategies. We examine the impact of different noise level selection strategies on SSL performance in Tab. 7. Three strategies are compared: Constant Selection (CS), Random Selection (RS), and Adaptive Selection (AS). CS applies a uniform noise level $c$ to all samples, with experiments conducted at various fixed levels (CS(0), CS(100), CS(200), CS(300), CS(400)). RS introduces variability by randomly selecting noise levels from the set ${0,100,200,300,400}$. AS dynamically adjusts noise levels based on the input image’s foreground proportion, as guided by Eq. (8). We use the same pretraining and linear evaluation settings as Tab. 6. The results indicate that AS achieves the highest accuracy at 73.96%, demonstrating the advantage of dynamically adjusting noise levels according to input characteristics. CS and RS yield lower performance, because static or random noise levels may result in overly similar or false positive pairs.
@@
Influence of the probability to apply GenView. The impact of different probabilities ($\alpha$) for applying GenView augmentation is shown in Tab. 8. An increase of the probability ($\alpha$) of applying GenView results in improved model performance, with top-1 accuracy consistently increasing from 62.39% at $\alpha=0$ to 70.55% at $\alpha=1.0$. This highlights the significance of a higher GenView application probability in enhancing the model’s ability to learn meaningful representations. By default, we set $\alpha=1$ for all the experiments in our main results.
@@ 4.3 Qualitative Evaluation
A qualitative illustration of the positive views constructed by GenView is shown in Fig. 4. The top rows display original images, and the bottom rows show images generated by GenView. This visualization demonstrates GenView’s capacity to introduce variations in background, pose, and view angle while preserving the main semantics, which is crucial for learning invariant representations. More visual examples are provided in the Appendix B.
@@ 5 Conclusion
Appendix A Implementation Details
@@
Method for adding noise perturbations. We use the pretrained Stable unCLIP v2-$1$https://huggingface.co/stabilityai/stable-diffusion-2-1-unclip model to generate image variations based on CLIP image embeddings $c$. An empty string serves as the text prompt to avoid any reference to image contexts or object names. The noised image embedding with perturbation strength $l$ is defined as: $\text{noisy}(\textbf{c},l)=\sqrt{\overline{a}_{l}}\mathbf{c}+\sqrt{(1-\overline{a}_{l})}\varepsilon$, where $\varepsilon\sim\mathcal{N}(0,\mathbf{I})$, and $\overline{a}_{l}$ is the cumulative product of $\alpha_{i}$ values for $i$ ranging from 0 to $l$. Each $\alpha_{i}$ is defined as $1-\beta_{i}$, with $\beta_{i}$ representing the noise variance introduced at step $i$, following the default linear schedule for $\beta_{[1:l]}$ from DDPM [37]. Higher values of the perturbation strength $l$ result in increased diversity in the generated images.
@@
Method for calculating foreground proportion. We use the pretrained CLIP ViT-H/14 backbone [^2] https://huggingface.co/laion/CLIP-ViT-H-14-laion2B-s32B-b79K, which serves as the conditional image encoder in Stable UnCLIP v2-1, for the encoder $C$ used in determining the proportion of foreground content before image generation. This backbone generates 256 tokens with a dimension of 1280 from a $224^{2}$ input resolution. For calculating PCA features, 10,000 images are randomly sampled from the original dataset. The threshold $a$ in Eq. 7 is selected to ensure that foreground tokens account for approximately 40% of the total tokens, providing a clear separation between foreground and background as depicted in the  Sec. 3.3.
@@
Method for generating attention maps. We employ the pretrained CLIP ConvNext-Base (with wide embedding dimension) backbone [^3] https://huggingface.co/laion/CLIP-convnext_base_w-laion2B-s13B-b82K-augreg as the encoder $E$ to extract feature maps from augmented positive views. These feature maps have a resolution of $7^{2}$ based on a $224^{2}$ input resolution. We compute foreground and background attention maps using the PCA vector computation method described in the previous paragraph.
@@
Hyper-parameters for view generation. We generate one augmented image for each image in the training set of IN-1K/CF10/CF100/TinyIN, with $T$ (the number of denoising steps) set to 20 for efficiency. The classifier-free guidance scale [36] is set to 10 to ensure image quality. The diversity of generated images is controlled by the level of noise perturbations applied to the image embeddings. To match the original dataset sizes of IN-1K/CF10/CF100/TinyIN, we resize the generated images from their original resolution of $768^{2}$ to $512^{2}$/$32^{2}$/$32^{2}$/$64^{2}$, respectively.
Comparison with SSL methods. The baseline results of MoCov3 in Tab. 1 are from the public codebase [^4] https://github.com/facebookresearch/moco-v3. Hyper-parameters for comparison with other SSL methods pretrained on IN-1K are listed in Tab. 9.
@@
Comparison with naive augmentation methods. To expand IN-1K with additional training data without introducing new classes, we employ a retrieval-based technique [5] for expanding IN-1K with Laion400M. We query the entire Laion400M dataset with 0.3 million randomly sampled IN-1K images and select the most similar image for each query image. For expanding IN-1K with IN-21K, we randomly sample 0.3 million non-repeating images with labels matching those in the IN-1K dataset. For GenView, positive views are generated for 0.15 million randomly sampled IN-1K images. For the experiments, the ResNet-50 models are pretrained on the expanded dataset using a batch size of 512. We apply a cosine decay learning rate schedule and use the LARS optimizer with a learning rate of 1.2, weight decay of 1e-6, and momentum of 0.9. Linear evaluation settings align with those in Tab. 1.

### v2

@@ Abstract
Keywords:
Self-supervised learning Contrastive learning View generation Generative models
footnotetext: * Corresponding authors
@@ 1 Introduction
Self-supervised learning (SSL) has demonstrated remarkable capability in acquiring robust and generalized visual representations from abundant unlabeled data sources [63, 27, 94, 64, 33, 16, 30, 106, 39, 32, 50, 2, 26, 90, 97, 15, 104], which can be transferred or leveraged in downstream tasks. Among the various approaches within SSL, Contrastive Learning (CL) [16, 17, 12, 13, 19, 106, 39, 52] has emerged as a prominent method, showcasing its effectiveness in numerous downstream tasks (e.g., classification [35, 51, 98], detection [72, 53, 34, 9, 87], and segmentation [28, 59, 92, 48, 49]). CL aims to learn invariant representations that remain consistent across various conditions or environments by maximizing the similarity of representations obtained from different distorted versions of a sample, referred to as positive views. Consequently, the construction of high-quality positive views is crucial for CL. A high-quality positive view should retain the semantics of the original images while introducing as much semantic-irrelevant attribute diversity and environmental variations as possible, such that the learned representations can be more generalizable for downstream tasks.
@@ 2 Related Work
Self-Supervised Learning.
@@
Instead of training generative models from scratch, some methods use pretrained generative models to augment representation learning, leveraging the prior knowledge learned from large-scale datasets [77, 14] to enhance the high-level diversity of the generated views [41, 36, 82, 22, 80, 85, 103]. However, these models rely on constant [36, 22, 82, 103] or random [80, 85, 107] hyperparameters to determine the extent of deviation in the generated images. This can lead to uncontrolled data generation characterized by inconsistent semantics with the conditional image, reducing the quality of positive pairs. In contrast, our approach employs adaptive view generation that controls the noise level when sampling images to keep a balance between semantic fidelity and diversity based on individual image characteristics. We also propose a quality-driven contrastive loss to enhance the contributions of high-quality positive pairs while diminishing the impact of low-quality and false pairs.
@@ 3.3 Adaptive View Generation
This process works in an offline manner before SSL training, so does not increase the burden on training time. Besides, the offline view generation is once-for-all and the generation result can be re-used multiple times for various baselines.
@@ 3.4 Quality-driven Contrastive Loss
$\mathbf{z}^{b}_{i}$ $=\mathbf{M}_{i}^{b}\otimes\mathbf{F}_{i},\quad\mathbf{z}^{b+}_{i}=\mathbf{M}_{i}^{b+}\otimes\mathbf{F}_{i}^{+},$
@@ 4 Experiments
We compare GenView with state-of-the-art SSL methods, including MoCov2 [33], BYOL [30], SwAV [12], SimSiam [18], and MoCov3 [19]. We experiment with various network architectures, such as ResNet-18 [35], ResNet-50 [35], ViT-S [21], and ViT-B [21]. By default, ResNet-50 serves as the backbone. ViT-S and ViT-B are adopted for comparison with MoCov3. For details on adaptive view generation and quality-driven contrastive loss implementations for different pretraining datasets, please refer to the Appendices 0.A and 0.C.
@@ 4.1 Main Results
Linear classification.
Table 1: Linear evaluation on IN-1K. *: our reproduction.
@@
GenView is framework-agnostic, allowing flexibility with SSL frameworks and associated training components like backbone networks, loss functions, and optimizers. To ensure fair comparisons, we maintain consistent pretraining settings as baseline methods on ImageNet-1K [20] (IN-1K). To evaluate our method, we follow a standard linear classification protocol, as described in previous works [16, 17, 30]. The linear classifier is trained on top of the frozen representation for 90 epochs with a batch size of 1,024, an initial learning rate of 0.4, an SGD optimizer with 0.9 momentum and no weight decay, and the cosine-annealed learning rate schedule [60]. For ViT-based models, the initial learning rate is set to 12. LABEL:tab:exp_linear presents the results of top-1 accuracy on the validation set of IN-1K. GenView consistently improves SSL performance across various frameworks, including ResNet-50 and Transformer architectures like ViT-S and ViT-B. Its effectiveness is maintained across different pretraining epochs, outperforming the MoCov3 baselines pretrained for 100 or 300 epochs. GenView outperforms C-Crop [66] that also constructs better views, highlighting our advantage in utilizing pretrained generative models’ prior knowledge to create diverse views in a controlled manner. GenView can complement both contrastive (e.g. MoCov2 and MoCov3) and non-contrastive methods (e.g. BYOL, SimSiam, and SwAV), addressing their limitations of positive pair quality. These results demonstrate GenView’s consistent ability in enhancing the linear classification performance of various SSL models. It’s noted that when GenView is integrated with MoCov3 utilizing a ResNet-50 backbone and pretrained over 300 epochs, it achieves competitive performance (74.8% with 1.28 million images) compared to CLIP (74.3% on WebImageText with 400 million pairs), highlighting GenView’s efficiency.
Table 2: Comparison with existing SSL methods for semi-supervised learning on IN-1K. Models with ResNet-50 backbone are pretrained on IN-1K. *: our reproduction.
@@
Table 3: Transfer learning on MS-COCO object detection and instance segmentation. Models with ResNet-50 backbone are pretrained for 200 epochs on IN-1K. *: our reproduction.
@@
Semi-supervised classification.
We evaluate the fine-tuning performance of the pretraind models for semi-supervised classification with 1% and 10% of labeled IN-1K samples, selected by SimCLR [16]. We fine-tune the models for 20 epochs with the classifier learning rate 1.0 (0.2) and backbone learning rate 0.00001 (0.02) for 1% (10%) subset with a cosine-annealed scheduler. Tab. 3 presents the results of top-1 and top-5 accuracy on the validation set of IN-1K. Our method consistently outperforms the baseline approaches across different training durations. With 1% labels, GenView pretrained for 200 epochs with MoCov2 achieves an improvement of +8.5% in top-1 accuracy, and the one pretrained for 300 epochs with MoCov3 still improves top-1 accuracy by +1.9%.
Transfer learning on object detection and instance segmentation.
@@
BYOL + GenView 93.56 67.53 54.79
Comparison with naive augmentation methods.
Comparison with other view construction methods.
@@ 4.2 Ablations
$\checkmark$ $\checkmark$ $\checkmark$ 75.40 ($\uparrow$ 9.88)
@@
Influence of each component.
We evaluate the contributions of individual components as well as their combinations. ResNet-18 models are pretrained on IN-100 for 100 epochs using MoCov3 as the baseline, with a batch size of 512. IN-100 is a subset of IN-1K selected by [83]. For conditioning the generation of positive views with GenView, we employ 50,000 randomly selected class-balanced images from IN-100. We use a cosine decay learning rate schedule and employ the LARS optimizer with a learning rate of 1.2, weight decay of 1e-6, and momentum of 0.9. Linear evaluation settings are consistent with those detailed in LABEL:tab:exp_linear, with a training duration of 50 epochs. Tab. 8 offers valuable insights: (1) Utilizing our framework but without our adaptive view generation significantly enhances accuracy, achieving a top-1 accuracy improvement of 5.98% compared to the baseline. (2) The incorporation of adaptive view generation further elevates model performance, resulting in an improvement of 8.44% (from 65.52% to 73.96%). (3) The quality-driven contrastive loss also plays a pivotal role in our framework. It can further improve the performance of adaptive view generation. Applying the quality-driven contrastive loss to the baseline method leads to a modest gain of 1.45% (from 65.52% to 66.97%). However, when combined with our framework, a more substantial performance improvement of 3.38% (from 71.50% to 74.88%) is observed. This highlights the effectiveness of our framework and also the importance of the proposed modules in enhancing contrastive learning by improving the quality of positive pairs.
Influence of the noise level selection strategies.
We examine the impact of different noise level selection strategies on SSL performance in Tab. 8. Three strategies are compared: Constant Selection (CS), Random Selection (RS), and Adaptive Selection (AS). CS applies a uniform noise level $c$ to all samples, with experiments conducted at various fixed levels (CS(0), CS(100), CS(200), CS(300), CS(400)). RS introduces variability by randomly selecting noise levels from the set ${0,100,200,300,400}$. AS dynamically adjusts noise levels based on the input image’s foreground proportion, as guided by Eq. (8). We use the same pretraining and linear evaluation settings as Tab. 8. The results indicate that AS achieves the highest accuracy at 73.96%, demonstrating the advantage of dynamically adjusting noise levels according to input characteristics. CS and RS yield lower performance, because static or random noise levels may result in overly similar or false positive pairs.
Influence of the probability to apply GenView.
The impact of different probabilities ($\alpha$) for applying GenView augmentation is shown in Tab. 8. An increase of the probability ($\alpha$) of applying GenView results in improved model performance, with top-1 accuracy consistently increasing from 62.39% at $\alpha=0$ to 70.55% at $\alpha=1.0$. This highlights the significance of a higher GenView application probability in enhancing the model’s ability to learn meaningful representations. By default, we set $\alpha=1$ for all the experiments in our main results.
@@ 4.3 Qualitative Evaluation
A qualitative illustration of the positive views constructed by GenView is shown in Fig. 4. The top rows display original images, and the bottom rows show images generated by GenView. This visualization demonstrates GenView’s capacity to introduce variations in background, pose, and view angle while preserving the main semantics, which is crucial for learning invariant representations. More visual examples are provided in the Appendix 0.B.
@@ 5 Conclusion
Acknowledgements
This work was supported in part by the National Natural Science Foundation of China under Grant 62376069, in part by Young Elite Scientists Sponsorship Program by CAST under Grant 2023QNRC001, and in part by Guangdong Basic and Applied Basic Research Foundation under Grant 2024A1515012027. The work was also supported by funding from KAUST Center of Excellence on GenAI, under award number 5940.
Appendix 0.A Implementation Details
Method for adding noise perturbations.
@@ Appendix 0.A Implementation Details
We use the pretrained Stable unCLIP v2-1https://huggingface.co/stabilityai/stable-diffusion-2-1-unclip model to generate image variations based on CLIP image embeddings $c$. An empty string serves as the text prompt to avoid any reference to image contexts or object names. The noised image embedding with perturbation strength $l$ is defined as: $\text{noisy}(\textbf{c},l)=\sqrt{\overline{a}_{l}}\mathbf{c}+\sqrt{(1-\overline{a}_{l})}\varepsilon$, where $\varepsilon\sim\mathcal{N}(0,\mathbf{I})$, and $\overline{a}_{l}$ is the cumulative product of $\alpha_{i}$ values for $i$ ranging from 0 to $l$. Each $\alpha_{i}$ is defined as $1-\beta_{i}$, with $\beta_{i}$ representing the noise variance introduced at step $i$, following the default linear schedule for $\beta_{[1:l]}$ from DDPM [37]. Higher values of the perturbation strength $l$ result in increased diversity in the generated images.
Method for calculating foreground proportion.
@@
We use the pretrained CLIP ViT-H/14 backbonehttps://huggingface.co/laion/CLIP-ViT-H-14-laion2B-s32B-b79K, which serves as the conditional image encoder in Stable UnCLIP v2-1, for the encoder $C$ used in determining the proportion of foreground content before image generation. This backbone generates 256 tokens with a dimension of 1280 from a $224^{2}$ input resolution. For calculating PCA features, 10,000 images are randomly sampled from the original dataset. The threshold $a$ in Eq. 7 is selected to ensure that foreground tokens account for approximately 40% of the total tokens, providing a clear separation between foreground and background as depicted in the  Sec. 3.3.
Method for generating attention maps.
@@
We employ the pretrained CLIP ConvNext-Base (with wide embedding dimension) backbonehttps://huggingface.co/laion/CLIP-convnext_base_w-laion2B-s13B-b82K-augreg as the encoder $E$ to extract feature maps from augmented positive views. These feature maps have a resolution of $7^{2}$ based on a $224^{2}$ input resolution. We compute foreground and background attention maps using the PCA vector computation method described in the previous paragraph.
Hyper-parameters for view generation.
@@
We generate one augmented image for each image in the training set of IN-1K/CF10/CF100/TinyIN, with $T$ (the number of denoising steps) set to 20 for efficiency. The classifier-free guidance scale [38] is set to 10 to ensure image quality. The diversity of generated images is controlled by the level of noise perturbations applied to the image embeddings. To match the original dataset sizes of IN-1K/CF10/CF100/TinyIN, we resize the generated images from their original resolution of $768^{2}$ to $512^{2}$/$32^{2}$/$32^{2}$/$64^{2}$, respectively.
Comparison with SSL methods.
@@
The baseline results of MoCov3 in LABEL:tab:exp_linear are from the public codebasehttps://github.com/facebookresearch/moco-v3. Hyper-parameters for comparison with other SSL methods pretrained on IN-1K are listed in Tab. 9.
Comparison with naive augmentation methods.
To expand IN-1K with additional training data without introducing new classes, we employ a retrieval-based technique [5] for expanding IN-1K with Laion400M. We query the entire Laion400M dataset with 0.3 million randomly sampled IN-1K images and select the most similar image for each query image. For expanding IN-1K with IN-21K, we randomly sample 0.3 million non-repeating images with labels matching those in the IN-1K dataset. For GenView, positive views are generated for 0.15 million randomly sampled IN-1K images. For the experiments, the ResNet-50 models are pretrained on the expanded dataset using a batch size of 512. We apply a cosine decay learning rate schedule and use the LARS optimizer with a learning rate of 1.2, weight decay of 1e-6, and momentum of 0.9. Linear evaluation settings align with those in LABEL:tab:exp_linear.
Comparison with other view construction methods.
We compare our approach with several baseline methods, including ContrastiveCrop [66] (C-Crop), ViewMaker [81], Neural Transform Network [43] (NTN), Local Manifold Augmentation method [99] (LMA), Diffusion-based augmentation from scratch [101] (DiffAug), and $\mathcal{W}$-perturb [31]. For our experiments, we use three datasets: CF10, CF100, and TinyIN. We employ SGD as the optimizer with a learning rate of 0.5, weight decay of 5e-4, and momentum of 0.9. The learning rate follows a linear warm-up for 10 epochs and then switches to the cosine decay scheduler. Batch sizes are set to 512 for CF10 and CF100 and 256 for TinyIN. The momentum coefficient for the momentum-updated encoder and memory buffer size is set to 0.99/0.99/0.996 and 4096/4096/16384 for CF10/CF100/TinyIN, respectively. We use the $\mathcal{L}_{\text{SSL, NCE}}$ loss with a temperature of 0.2 and train for 500 epochs on CF10 and CF100 and 200 epochs on TinyIN. The backbone architecture used is ResNet18 with an embedding dimension of 512 and a projection dimension of 128. We replace the first 7x7 Conv of stride 2 with 3x3 Conv of stride 1 and remove the first max-pooling operation. For data augmentations, we use random resized crops (the lower bound of random crop ratio is set to 0.2), color distortion (strength=0.4) with a probability of 0.8, and Gaussian blur with a probability of 0.5. The images from the CF10/CF100 and TinyIN datasets are resized to 32x32 and 64x64 resolution.
Appendix 0.B Additional Illustration
Further visualization and limitation analysis.